OpenStudy (anonymous):

If the point (4,-1) is a point on the graph of f, the f(_)=_

jimthompson5910 (jim_thompson5910):

(4,-1) means x = 4 and y = -1 since every point is of the form (x,y)

jimthompson5910 (jim_thompson5910):

it turns out that y = f(x) so if x = 4 and y = -1 then y = f(x) -1 = f(4) ... plug in x = 4 and y = -1 f(4) = -1

jimthompson5910 (jim_thompson5910):

x is the input of the function y is the output of the function

OpenStudy (anonymous):

Thanks that makes a lot more sense now

jimthompson5910 (jim_thompson5910):

you're welcome